OpenStudy (purplevampireeclipse):

Replace ? with =, >, or < to make the statement true. 18 ÷ 6 + 3 ? 6 + 12 ÷ 3 A. = B. > C. <

OpenStudy (purplevampireeclipse):

a?

OpenStudy (goku_1999):

C

OpenStudy (3mar):

Let's think together! What is the most priority? */ or +-?

OpenStudy (purplevampireeclipse):

divide and addition

OpenStudy (3mar):

Good. Let's apply that for the above ...

OpenStudy (3mar):

you mean divide and multiplication?

OpenStudy (purplevampireeclipse):

no

OpenStudy (3mar):

\[18÷6+3=(18÷6)+3=3+3=1\]ok?

OpenStudy (purplevampireeclipse):

okay

OpenStudy (3mar):

the other side: \[6+12÷3=6+(12÷3)=6+4=10\]follow?

OpenStudy (purplevampireeclipse):

yes

OpenStudy (3mar):

sorryyyy \[18÷6+3=(18÷6)+3=3+3=6\]

OpenStudy (3mar):

so what is the relation between 6 and 10?

OpenStudy (3mar):

You pick the answer while I am praying. Salam!

OpenStudy (3mar):

Thank you for the medal!